OpenStudy (anonymous):

Show that the equation 2^x = 2 − x^2 has at least two real roots

OpenStudy (anonymous):

Solve for 0, then factor.

OpenStudy (amistre64):

use the determinate..i think its called of the quad formula.... if its 0 only 1 root, if it (+) 2 roots, it its (-) imaginary roots

OpenStudy (anonymous):

but its 2^x an exponential fuction...so how to use quadratic here ?

OpenStudy (amistre64):

thought that was a typo lol

myininaya (myininaya):

its called the discriminant

OpenStudy (anonymous):

Same here.

OpenStudy (amistre64):

soooo close......

myininaya (myininaya):

b^2-4ac

OpenStudy (anonymous):

but how does discriminant help ?

OpenStudy (anonymous):

\[-x^2+2^x -2 = 0\]

OpenStudy (amistre64):

x^2 = -2^x +2 square root both sides?

OpenStudy (amistre64):

x = +-sqrt(-2^x +2)

OpenStudy (anonymous):

sry but i m still in a fix ....jeezzzzz

OpenStudy (amistre64):

that doesnt tell you the roots, but it might show that there are 2 of them :)

OpenStudy (anonymous):

ahh..got ya !!!

OpenStudy (anonymous):

hey but the roots can be complex by taking the roots???

OpenStudy (anonymous):

by taking the square root

OpenStudy (amistre64):

roots can be complex and most are... but the only one to concern yourself with are the "real" roots

OpenStudy (anonymous):

yes...when it was +- sqrt( 2- 2^x). that does not show that v r gonna get two real roots which is the question asking for?

OpenStudy (amistre64):

if you wanna limit your self to even values of x

OpenStudy (anonymous):

but there is no domain given in the question

OpenStudy (anonymous):

meanwhile i checked on wolfram alpha

OpenStudy (amistre64):

oh wait... is it (-2)^x or -(2^x)??

OpenStudy (anonymous):

they had a graph which cuts on two points

OpenStudy (anonymous):

the latter one

OpenStudy (amistre64):

maybe log rules would help

OpenStudy (anonymous):

would you know guys what is uniform continuity?

OpenStudy (amistre64):

log(2^x) = log(2-x^2) x log(2) = log(2-x^2) log(2) = log(2-x^2)^1/x 2 =xroot(2-x^2) ??

OpenStudy (amistre64):

uniform?... id have to look it up

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f(x) = sin x is continuous on R. Is it uniformly continuous on R?

OpenStudy (amistre64):

id vote yes, its periodic, repetitive, and constantly the "same"

OpenStudy (amistre64):

if "x" is a fraction that gets us back to integer exponents with the first question....

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2^x+%3D+2+%E2%88%92+x^2 check out that one for me plsssssss

OpenStudy (amistre64):

..... yeah.....thats correct :)

OpenStudy (anonymous):

so u think graph would be enough?

OpenStudy (amistre64):

the graph itself is visual proof, the mathematical proof seems a but difficult to reach.

OpenStudy (anonymous):

yes...and the second question like you said or is there any better way?

OpenStudy (amistre64):

if you graph the sin function, it sounds as if it would fit a "uniform" continuity.

OpenStudy (anonymous):

yes...thank you.... can i ask you anything else or i m just pissing u off :(

OpenStudy (amistre64):

lol.... you can ask....i might not have an answer tho :)

OpenStudy (anonymous):

how to write pi as a symbol on computer?

OpenStudy (amistre64):

on any computer? or in here?

OpenStudy (anonymous):

any computer

OpenStudy (amistre64):

its the ascii code 227 i think, there should also be a character keyboard in your menu that will bring up a virtual keyboard to find it on

OpenStudy (anonymous):

yeah mayb in microsoft word ...they might have it ...thanx

OpenStudy (amistre64):

programs..accessories....system tools....character maap is where mine is

OpenStudy (anonymous):

hey let me check that one ...thanx

OpenStudy (amistre64):

П lol

OpenStudy (anonymous):

hey u got that...i might copy it n save it on word ...lol

OpenStudy (anonymous):

ППППППППППП hhahah

OpenStudy (amistre64):

Пп hehehe

OpenStudy (anonymous):

lim x->pi/3 ((3+2cosx)^1/2 - 2) / ln (1+sin3x) ??? whats ur views on that one?

OpenStudy (amistre64):

pi/3 = 60 degrees..... is [sqrt(3+2cos(x)) - 2] / ln(1+sin(3x)) right?

OpenStudy (anonymous):

exactly

OpenStudy (amistre64):

sqrt(3+2cos(60)) - 2 sqrt(3 + 2(1/2)) - 2 ------------------- = ----------------- ln(1+sin(180)) ln(1+0)

OpenStudy (anonymous):

but thats a limit question ...plugging in the values making it 0/0

OpenStudy (amistre64):

sqrt(4) - 2 0 --------- = --- ln(1) 0 I see the dellima

OpenStudy (amistre64):

the derivative of each would be....

OpenStudy (amistre64):

Dx(ln(x)) = 1/x Dx(ln(1+sin(3x))) = 3cos(3x)/1+sin(3x) the reciprocal of that is what youll need

OpenStudy (amistre64):

Dx(sqrt(3+2cos(x))-2) = -sin(x)/sqrt(3+2cos(x))

OpenStudy (amistre64):

-sin(x) (1+sin(3x)) ---------------------- sqrt(3+2cos(x)) (3cos(3x)

OpenStudy (anonymous):

yes i was also somewhere there

OpenStudy (amistre64):

-sqrt(3) sqrt(3) --------- = -------- is what I get 2(3)(-1)(2) 12

OpenStudy (anonymous):

i have got 1/4sqrt3

OpenStudy (amistre64):

same thing, rationalize your denominator

OpenStudy (amistre64):

sqrt(3)/4(3)

OpenStudy (anonymous):

got ya !!! u must b a teacher for sure ... aren't u ?

OpenStudy (amistre64):

lol ..... nope, just an unemployed carpenter

OpenStudy (anonymous):

wtf??? impossible ...to me u r a GREAT TEACHER !!!! and more of a great human being

OpenStudy (amistre64):

thanx :)

OpenStudy (anonymous):

u have been so helpful...these questions were all for my next week homwork assignment...now i m so relaxed...thanx bro....really appreciate ur help

OpenStudy (anonymous):

Refer to the attachment.